«Homework #2 Problem 19.27 Review Part A A container holds 0.800 g of oxygen at a pressure of 8.80 atm How much heat is required to increase the temperature by 118 °C at constant pressure? Express your answer with the appropriate units. 124.18 Submit Incorrect Try Again; 4 attempts remaining Part B How much will the temperature increase if this amount of heat energy is transferred to the gas at constant volume? Express your answer with the appropriate units. oC Submit Reguest Answer Provide Feedback
Answer #1

part A)

specific heat at constant pressure , Cp = 3.5 * R

Cp = 3.5 * 8.314 = 29.1 J/mol

heat needed = 0.80/32 * 29.1 * 118

heat needed = 85.8 J

part B) at constant volume

Cv = 2.5 * R = 2.5 * 8.314 = 20.8 J/mol.degree C

let the temperature change is del T

85.8 = 0.80/32 * 20.8 * del T

del T = 165 degree C

the increase in temperature at constant volume is 165 dgeree